# Efficient Inference in First Passage Time Models

### v1

@@ 5 Numerical Examples
In this section, we provide several numerical examples to demonstrate the effectiveness of our method. In Examples 5.1, 5.2, and 5.3, we compute exact or approximate FPTDs using Algorithm 1 and verify the results by comparing them with simulated data. In Section 5.4, we showcase the speed of Algorithm 1 for likelihood computation and statistical inference.
@@
$\widehat{X}(r_{n+1})=\widehat{X}(r_{n})+\mu(\widehat{X}(r_{n}),r_{n})(r_{n+1}-r_{n})+\sigma(\widehat{X}(r_{n}),r_{n})\Delta W(r_{n})$
to simulate an approximate sample path of $X$, where the random variable $\Delta W(r_{n})=W(r_{n+1})-W(r_{n})$ is distributed as $\mathcal{N}(0,r_{n+1}-r_{n})$, and all $\Delta W(r_{n})$’s are independent. Such simulation will of course be biased due to the unavoidable discretization error. Then, the first passage time data is the first time $r_{n}$ such that $\widehat{X}(r_{n})\notin(\ell(r_{n}),u(r_{n}))$. However, even if we could simulate $X$ exactly on the discrete grid $\{r_{n}\}_{n=1}^{\infty}$, this would not sample $\tau$ exactly, as it is possible for $X$ to cross the barrier between grid points $r_{n}$ and $r_{n+1}$ without exceeding the barrier at either of these grid points.
@@ 5.1 FPTDs for GDDMs with piecewise Constant Drift
starting at $x_{0}=-0.5$. This is a generic example of a multi-stage model. Unlike the aDDM, here the breakpoints of the drift function are the same for all trials.
We use our Algorithm 1 to compute the FPTDs on the upper boundary and lower boundary and compare them with the normalized histogram obtained from 50000 Monte Carlo samples. The results are shown in Figure 5. The FPTD exhibits corners at all breakpoints of $\mu(t)$. Despite this complexity, our result visually matches the empirical densities.
@@ 5.2 FPTDs for GDDMs with Nonlinear Drift and Boundaries
$\widetilde{u}(t)=\frac{1}{\sigma}\Big{(}u(t)-\int_{0}^{t}\mu(r)\,\mathrm{d}r\Big{)},\quad\widetilde{\ell}(t)=\frac{1}{\sigma}\Big{(}\ell(t)-\int_{0}^{t}\mu(r)\,\mathrm{d}r\Big{)}$
and work with the transformed model $\widetilde{X}(t)=\frac{X(0)}{\sigma}+W(t)$. We approximate $\widetilde{u}(t)$ and ~$\widetilde{\ell}(t)$ by piecewise linear functions and use Algorithm 1 to evaluate the FPTD of the approximate multi-stage model. This can apply to any nonlinear drift and boundaries.
@@
starting at $x_{0}=-0.5$. The upper boundary $u(t)$ is the survival function of a Weibull$(5,3)$ distribution, and $\ell(t)$ is its reflection about the $t$-axis. The particular choice of a Weibull distribution follows relevant examples from prior work [18].
The original boundaries, and the transformed boundaries with their piecewise linear approximations are shown in Figure 6 and Figure 6, respectively. The computed approximate FPTDs and the histogram consisting of 50,000 samples are displayed in Figure 6. The approximate FPTD and the histogram agree well.
(a) Original GDDM.
(b) Transformed GDDM.
(c) Plots of FPTDs and histogram.
Figure 6: Example 5.2: a GDDM with sinusoidal drift and Weibull survival boundaries. (6) and (6) illustrate the GDDM before and after transformation. The expected value of the process (blue curves) is transformed from a sinusoidal function to a constant zero. In (6), red dots on the upper boundary and purple dots on the lower boundary indicate selected interpolation points for approximating the respective boundaries as piecewise linear functions. (6) displays first passage times: those on the upper boundary appear on the positive $t$-axis, while those on the lower boundary appear on the negative $t$-axis, forming a valid probability density function.
@@ 5.3 FPTDs for Ornstein-Uhlenbeck Models
$\widetilde{u}(s)$ $=\frac{1}{\sigma}\Big{(}u\Big{(}\frac{1}{2\theta}\log(1+2\theta s)\Big{)}\sqrt{1+2\theta s}-\lambda\sqrt{1+2\theta s}+\lambda\Big{)}$
~$\widetilde{\ell}(s)$ $=\frac{1}{\sigma}\Big{(}\ell\Big{(}\frac{1}{2\theta}\log(1+2\theta s)\Big{)}\sqrt{1+2\theta s}-\lambda\sqrt{1+2\theta s}+\lambda\Big{)}$
@@
5.4 Statistical Computing for aDDMs
In the following examples, we show the effectiveness and efficiency of our method for computing the likelihood and inferring parameters for aDDMs. The aDDM incorporates visual attention as a key factor underlying the decision process behind simple choice behavior. It adjusts the drift rate $\mu(t)$ based on instantaneous attention within a given experiment trial, adopting values $\mu_{1}$ or $\mu_{2}$ depending on whether the participant fixates on choice options A or B at time $t$. Specifically, if we let $V(t)$ be the process taking values in $\{\mathrm{A},\mathrm{B}\}$ representing the participant’s visual fixation at time $t$, then the aDDM corresponds to the following SDE (conditioned on $V$),
$\,\mathrm{d}X(t)=A(V(t))\,\mathrm{d}t+\sigma\,\mathrm{d}W(t),$ (20)
where $A(v)=\begin{cases}\mu_{1}&\text{if }v=\mathrm{A},\\ \mu_{2}&\text{if }v=\mathrm{B}.\\ \end{cases}$
Note that $\mu(t)=A(V(t))$ is a piecewise constant function as in the multistage model of Section 4.3, but that it will generally be different on each experimental trial, since a participant’s visual fixation process will generally be different on each trial.
In an experimental setting, a participant’s fixation trajectory $V(t)$ is captured using an eye-tracker and can be viewed as a known covariate. Because the fixations drive time-varying drift functions, the configuration $\boldsymbol{\Xi}$ of the GDDM will be different on each experimental trial. Letting $\boldsymbol{\Xi}_{i}$ denote the configuration on the $i$-th trial and recalling that $\boldsymbol{\Xi}_{i}$ depends on unknown parameters $\boldsymbol{\theta}$, such as $\mu_{1}$ and $\mu_{2}$, that are shared by all trials and on the known visual fixation trajectory, say $V_{i}(t)$, for the subject’s $i$-th trial, the likelihood of $\boldsymbol{\theta}$ is
$\mathcal{L}_{\text{aDDM}}(\boldsymbol{\theta}\mid\{(\tau_{i},C_{i})\}_{i=1}^{n})=\prod_{i=1}^{n}f(\tau_{i},C_{i}\mid\boldsymbol{\Xi}_{i}).$ (21)
Here, we simulate a collection of $n=50000$ visual fixation trajectories $(V_{i})_{i=1}^{n}$ and the corresponding dataset $((\tau_{i},C_{i}))_{i=1}^{n}$ from a prescribed aDDM model with alternating drifts $\mu_{1},\mu_{2}$, constant diffusion $\sigma=1$ and symmetric linear collapsing boundaries
@@ 5.4 Statistical Computing for aDDMs
starting at $x_{0}\in(-a,a)$. The fixation trajectory $V_{i}(t)$ on the $i$-th trial was sampled from a gamma process to mimic the switching of attention in psychological experiments. The average number of fixations per trial in our simulation was 5.86.
@@ 5.4.1 Likelihoods for aDDM given a Dataset
We evaluate the performance of Algorithm 1 for computing likelihoods of the parameters given the dataset $((\tau_{i},C_{i}))_{i=1}^{n}$. The speed of our methods is compared with PyDDM [57], a prominent, widely applied library which serves as a benchmark on speed for the KFE approach (Section 3). When using PyDDM, we chose the finite difference discretization parameters ($\Delta t=0.001,\Delta x=0.001$) to be sufficiently small so that the single-trial FPTDs are visually indistinguishable from our solution.
Methods Time$(\mathrm{s})$
Algorithm 1 (1 thread) $5.113$
Algorithm 1 (8 threads) $0.682$
Algorithm 1 (32 threads) $0.275$
KFE (1 thread) $17561.83$
Table 2: Average times to evaluate the aDDM likelihood in equation (21): Algorithm 1 versus KFE as implemented in PyDDM. Times are average CPU times of $10$ repeated evaluations on an AMD EPYC 9684X Processor.
Table 2 shows that our method outperforms the KFE method by more than three orders of magnitude. Moreover, our method can easily be further accelerated with parallelization. The speedup of our method for aDDMs largely stems from its ability to progress in large time steps aligned with saccade times, leading to extremely fast single trial computation speed. In contrast, the KFE method requires a much smaller time discretization step to achieve similar accuracy. The accurate analytical formulas in equations (6)–(8) and the high-order efficient quadrature rules in equation (23) in the Appendix further contribute to the method’s speed as they are more accurate than the Crank-Nicolson scheme used in PyDDM.
@@ 5.4.2 Statistical Inference of aDDM Parameters
Here we conduct a parameter recovery study for the parameters $\boldsymbol{\theta}=(\mu_{1},\mu_{2},a,b,x_{0})$ using the maximum likelihood estimator (MLE):
$\widehat{\boldsymbol{\theta}}_{n}^{\text{\sc mle}}=\operatorname*{\arg\!\max}_{\boldsymbol{\theta}}\mathcal{L}_{\text{aDDM}}(\boldsymbol{\theta}\mid\{(\tau_{i},C_{i}\}_{i=1}^{n})=\operatorname*{\arg\!\max}_{\boldsymbol{\theta}}\prod_{i=1}^{n}f(\tau_{i},C_{i}\mid\boldsymbol{\Xi}_{i})$
We use Algorithm 1 to evaluate the likelihood and the trust region method for optimization. The true and estimated parameters are shown in Table 3 along with component-wise nonparametric pivotal bootstrap confidence intervals.
True Estimated 95% CI
$\mu_{1}$ $1.0$ $0.999$ $[0.985,1.015]$
$\mu_{2}$ $-0.8$ $-0.805$ $[-0.820,-0.790]$
$a$ $2.1$ $2.105$ $[2.093,2.116]$
$b$ $0.3$ $0.303$ $[0.298,0.306]$
$x_{0}$ $-0.2$ $-0.197$ $[-0.209,-0.186]$
Table 3: The true and estimated parameters of the aDDM with component-wise nonparametric pivotal bootstrap confidence intervals (based on 1000 bootstrap resamples). All the true parameters fall in their $95\%$ confidence intervals.
Numerical optimization using our method takes on the order of $1$ minute with $32$ threads, achievable even on a standard desktop computer. In contrast, as it takes almost $5$ hours to perform just a single likelihood evaluation using the KFE approach, the numerical optimization would take days or more likely weeks. This illustrates that our method enables fast and accurate parameter recovery.

### v2

@@ 5 Numerical Examples
In this section, we provide several numerical examples to demonstrate the effectiveness of our method. In Examples 5.1, 5.2, and 5.3, we compute exact or approximate FPTDs using Algorithm 1 and verify the results by comparing them with simulated data. In Section 5.4, we demonstrate the efficiency of Algorithm 1 for likelihood computation, highlighting its speed relative to previous methods and its applicability to statistical inference.
@@
$\widehat{X}(r_{n+1})=$ $\widehat{X}(r_{n})+\mu(\widehat{X}(r_{n}),r_{n})(r_{n+1}-r_{n})$
$+\sigma(\widehat{X}(r_{n}),r_{n})\Delta W(r_{n})$
to simulate an approximate sample path of $X$, where the random variable $\Delta W(r_{n})=W(r_{n+1})-W(r_{n})$ is distributed as $\mathcal{N}(0,r_{n+1}-r_{n})$, and all $\Delta W(r_{n})$’s are independent. Such a simulation will, of course, be biased due to the unavoidable discretization error. Then, the first passage time data is the first time $r_{n}$ such that $\widehat{X}(r_{n})\notin(\ell(r_{n}),u(r_{n}))$. However, even if we could simulate $X$ exactly on the discrete grid $(r_{n})_{n=1}^{\infty}$, this would not sample $\tau$ exactly, as it is possible for $X$ to cross the barrier between grid points $r_{n}$ and $r_{n+1}$ without exceeding the barrier at either of these grid points.
@@ 5.1 FPTDs for GDDMs with Piecewise Constant Drift
starting at $x_{0}=-0.5$. This is a generic example of a multi-stage model with the same drift terms for all trials.
We use our Algorithm 1 to compute the FPTDs on the upper boundary and lower boundary and compare them with the normalized histogram obtained from 50,000 Monte Carlo samples. The results are shown in Figure 5. The FPTD exhibits corners at all breakpoints of $\mu(t)$. Despite this complexity, our result visually shows excellent agreement with the empirical densities.
@@ 5.2 FPTDs for GDDMs with Nonlinear Drift and Boundaries
$\widetilde{u}(t)$ $=\frac{1}{\sigma}\Big(u(t)-\int_{0}^{t}\mu(r)\,\mathrm{d}r\Big),$
$\widetilde{\ell}(t)$ $=\frac{1}{\sigma}\Big(\ell(t)-\int_{0}^{t}\mu(r)\,\mathrm{d}r\Big)$
and work with the transformed model $\widetilde{X}(t)=\frac{X(0)}{\sigma}+W(t)$. We approximate $\widetilde{u}(t)$ and $\widetilde{\ell}(t)$ by piecewise linear functions and use Algorithm 1 to evaluate the FPTD of the approximate multi-stage model. This can apply to any nonlinear drift and boundaries.
@@
starting at $x_{0}=-0.5$. The upper boundary $u(t)$ is the survival function of a Weibull(5,3)(5,3) distribution, and $\ell(t)$ is its reflection about the $t$-axis. The choice of a Weibull distribution is guided by relevant examples in prior work [13, 19].
The original boundaries, and the transformed boundaries with their piecewise linear approximations are shown in Figure 6(a) and Figure 6(b), respectively. The computed approximate FPTDs and the histogram consisting of 50,000 samples are displayed in Figure 6(c). The approximate FPTD and the histogram agree well.
Figure 6: Example 5.2: a GDDM with sinusoidal drift and Weibull survival boundaries. (a) and (b) illustrate the GDDM before and after transformation. The expected value of the process (blue curves) is transformed from a sinusoidal function to a constant zero. In (b), red dots on the upper boundary and purple dots on the lower boundary indicate selected interpolation points for approximating the respective boundaries as piecewise linear functions. (c) displays first passage times: those on the upper boundary appear on the positive $t$-axis, while those on the lower boundary appear on the negative $t$-axis, forming a valid probability density function.
@@ 5.3 FPTDs for Ornstein-Uhlenbeck Models
$\widetilde{u}(s)=$ $\tfrac{1}{\sigma}\Big(u\big(\tfrac{1}{2\theta}\log(1+2\theta s)\big)\sqrt{1+2\theta s}$
$-\lambda\sqrt{1+2\theta s}+\lambda\Big),$
$\widetilde{\ell}(s)=$ $\tfrac{1}{\sigma}\Big(\ell\big(\tfrac{1}{2\theta}\log(1+2\theta s)\big)\sqrt{1+2\theta s}$
$-\lambda\sqrt{1+2\theta s}+\lambda\Big)$
@@
5.4 Practical Statistical Computing for aDDMs
In the following examples, we show the effectiveness and efficiency of our method for computing the likelihood and inferring parameters for aDDMs. The aDDM incorporates visual attention as a key factor underlying the decision process behind simple choice behavior. It adjusts the drift rate $\mu(t)$ based on instantaneous attention within a given experimental trial, adopting different values depending on whether the participant fixates on choice options A or B at time $t$. Let $V(t)$ be the process taking values in $\{\mathrm{A},\mathrm{B}\}$ representing the participant’s visual fixation at time $t$, and let $r_{A}$ and $r_{B}$ denote the participant’s numerical ratings of their preferences for options A and B, respectively. The aDDM corresponds to the following SDE (conditioned on $V,r_{A},r_{B}$) [29],
$\,\mathrm{d}X(t)=A(V(t),r_{A},r_{B})\,\mathrm{d}t+\sigma\,\mathrm{d}W(t)$ (20)
where
[…]
$\eta\in(0,1)$ is a parameter that reflects the discounting of the non-fixated option, and $\kappa>0$ parametrizes the overall magnitude of the drift. Note that $\mu(t)=A(V(t),r_{A},r_{B})$ is a piecewise constant function as in the multi-stage model of Section 4.3, but that it will generally be different on each experimental trial, both because a participant’s visual fixation process will generally be different on each trial and because the options presented to the participant can change across trials.
In an experimental setting, a participant’s fixation trajectory $V(t)$ is captured using an eye-tracker, while the participant’s preference ratings $r_{A}$ and $r_{B}$ are collected prior to the choice experiment; together, these can be viewed as known covariates. Because those covariates drive time-varying drift functions, the configuration $\boldsymbol{\Xi}$ of the GDDM will be different on each experimental trial. Letting $\boldsymbol{\Xi}_{i}$ denote the configuration on the $i$-th trial, then the likelihood of $\boldsymbol{\theta}$ is
$\mathcal{L}_{\text{aDDM}}(\boldsymbol{\theta}\mid((\tau_{i},C_{i}))_{i=1}^{n})=\prod_{i=1}^{n}f(\tau_{i},C_{i}\mid\boldsymbol{\Xi}_{i}).$ (21)
$\boldsymbol{\Xi}_{i}$ depends on unknown parameters $\boldsymbol{\theta}$ (e.g. $\eta$ and $\kappa$) that are shared by all trials, as well as on the trial-specific known visual fixation trajectory $V_{i}(t)$ and ratings $r_{A,i},r_{B,i}$ for the subject’s $i$-th trial.
Here, we simulate a collection of $n=50,000$ visual fixation trajectories $(V_{i}(t))_{i=1}^{n}$, preference ratings $((r_{A,i},r_{B,i}))_{i=1}^{n}$ and the corresponding dataset of choices and reaction times $((\tau_{i},C_{i}))_{i=1}^{n}$ from a prescribed aDDM model with drift parameters $\kappa,\eta$, constant diffusion $\sigma=1$ and symmetric linear collapsing boundaries
@@ 5.4 Practical Statistical Computing for aDDMs
starting at $x_{0}\in(-a,a)$. The fixation trajectory $V_{i}(t)$ on each trial was sampled from a gamma process to mimic the switching of attention in psychological experiments, and the ratings $r_{A,i},r_{B,i}$ are sampled from a discrete uniform distribution on $\{1,2,3,4,5\}$. The average number of fixations per trial in our simulation is 3.8.
@@ 5.4.1 Likelihoods for aDDM given a Dataset
We evaluate the performance of Algorithm 1 for computing likelihoods of the parameters at their true values given the dataset $((\tau_{i},C_{i}))_{i=1}^{n}$. We compare the speed of our methods with PyDDM [50], a prominent and highly optimized library which we use as a benchmark on speed for the KFE approach (Section 3). When using PyDDM, we can choose the finite difference discretization parameters $\Delta t$ and $\Delta x$ which determine the fineness of the temporal and spatial meshes, and as $\Delta t\rightarrow 0,\Delta x\rightarrow 0$, the numerical solution converges to the true solution of the underlying KFE governing the diffusion process. We test Algorithm 1 with varying quadrature orders and compare its speed–accuracy trade-off against the KFE method evaluated with different choices of $\Delta t$ and $\Delta x$. The results are shown in Figure 8.
Figure 8: Speed–accuracy trade-off for Algorithm 1 (blue) and the KFE method as implemented in PyDDM (red solid and dashed). Algorithm 1 performance is shown for increasing quadrature order from $10$ to $35$. KFE performance is shown for decreasing space-time discretization ($\Delta t=\Delta x$) from $0.05$ to $0.001$. In all cases, we used the corresponding method to compute the FPTDs for all trials: $(f(\tau_{i},C_{i}\mid\boldsymbol{\Xi}_{i}))_{i=1}^{n}$, where n=50,000n=50{,}000 and the parameters in $\boldsymbol{\Xi}_{i}$ are those that generated the data. CPU time in each case is the total time to compute FPTDs for all 50,000 trials. The FPTDs can be combined to give the trial-normalized negative log-likelihood $\ell^{*}=-\frac{1}{n}\sum_{i=1}^{n}\log f(\tau_{i},C_{i}\mid\boldsymbol{\Xi}_{i})$ (see Eq. (21)). Since the aDDM in this example corresponds exactly to a multi-stage GDDM, we used Algorithm 1 with a high quadrature order ($200$ quadrature points per stage) to obtain reference values for $\ell^{*}$ and $\ell^{\prime}$ (defined below). Relative errors for Algorithm 1 (blue line, vertical axis) are reported against this reference value of $\ell^{*}$. In four trials the KFE implementation (with these $\Delta t,\Delta x$) gave FPTDs of zero, yielding an infinite approximation of $\ell^{*}$. To provide a more meaningful sense of error for KFE, we also calculated the likelihood excluding those trials, namely $\ell^{\prime}=-\frac{1}{n-4}\sum_{\text{valid $i$}}\log f(\tau_{i},C_{i}\mid\boldsymbol{\Xi}_{i})$, where the “valid” $i$ are the 49,99649{,}996 trials for which KFE gave strictly positive FPTD approximations. The solid red line compares the KFE approximation of $\ell^{\prime}$ to the reference value of $\ell^{*}$. The error plateaus because KFE is forced to exclude the trials where its approximation fails. The dashed red line compares the KFE approximation of $\ell^{\prime}$ to the reference value of $\ell^{\prime}$. Although this is the wrong reference value for the target problem, it illustrates that the KFE approximation continues to improve as the discretization gets finer.
Figure 8 demonstrates that our method outperforms the KFE method by more than three orders of magnitude at a given target accuracy level. Moreover, our method is more numerically stable when evaluating trials with small likelihoods. PyDDM returned a likelihood of zero for several trials, complicating its use for likelihood-based methods.
The speedup of our method for aDDMs largely stems from its ability to progress in large time steps aligned with saccade times, leading to extremely fast single trial computation speed. The accurate analytical formulas in equations (6)–(8) and the high-order efficient quadrature rules in equation (24) further contribute to our method’s speed. In contrast, the KFE method requires much finer time discretization in the backward Euler scheme and therefore substantially more floating-point operations to attain comparable accuracy.
@@ 5.4.2 Statistical Inference of aDDM Parameters
Here we conduct parameter recovery studies for the parameters $\boldsymbol{\theta}=(\eta,\kappa,a,b,x_{0})$ in both the frequentist and Bayesian settings. We use the maximum likelihood estimator (MLE)
$\widehat{\boldsymbol{\theta}}_{n}^{\text{\sc mle}}$ $=\operatorname*{\arg\!\max}_{\boldsymbol{\theta}}\mathcal{L}_{\text{aDDM}}(\boldsymbol{\theta}\mid((\tau_{i},C_{i}))_{i=1}^{n})$
$=\operatorname*{\arg\!\max}_{\boldsymbol{\theta}}\prod_{i=1}^{n}f(\tau_{i},C_{i}\mid\boldsymbol{\Xi}_{i})$
as the frequentist approach. In the Bayesian framework, we place a prior $f_{\text{prior}}(\boldsymbol{\theta})$ on the parameters and approximate the posterior density
$f_{\text{posterior}}(\boldsymbol{\theta})\propto f_{\text{prior}}(\boldsymbol{\theta})\prod_{i=1}^{n}f(\tau_{i},C_{i}\mid\boldsymbol{\Xi}_{i})$
using Markov chain Monte Carlo (MCMC). For both approaches we use Algorithm 1 to evaluate the likelihood. The true and estimated parameters are shown in Table 2.
Table 2: True values, point estimates and interval estimates of aDDM parameters. The maximum likelihood estimates are reported together with component-wise nonparametric pivotal bootstrap confidence intervals based on 1000 resamples. The optimization is carried out using the trust region method. For the Bayesian approach, we report the posterior means and 95% Bayesian credible intervals. The MCMC sampling was performed using the Metropolis–Hastings algorithm with $10^{4}$ burn-in iterations followed by $10^{5}$ draws. The MLE, posterior means and medians (not reported) coincide up to three decimal places. Note that all true parameter values lie within their corresponding 95% confidence intervals and credible intervals.
True MLE 95% confidence interval posterior mean 95% Bayesian credible interval
$\eta$ $0.7$ $0.703$ $[0.698,0.708]$ $0.703$ $[0.697,0.709]$
$\kappa$ $0.5$ $0.506$ $[0.500,0.514]$ $0.506$ $[0.499,0.513]$
$a$ $2.1$ $2.109$ $[2.099,2.121]$ $2.109$ $[2.098,2.120]$
$b$ $0.3$ $0.302$ $[0.298,0.306]$ $0.302$ $[0.298,0.306]$
$x_{0}$ 0.2-0.2\ \ 0.195-0.195\ \ [0.206,0.184][-0.206,-0.184] 0.195-0.195\ \ [0.207,0.183][-0.207,-0.183]
For maximum likelihood estimation, the numerical optimization takes around 1 minute when run with 32 threads, a computation time that is readily achievable even on a standard desktop computer. For Bayesian inference, a MCMC run with $10^{5}$ draws completes in $10$ hours. In contrast, because a single round of accurate likelihood evaluation with the KFE approach alone requires nearly $4$ hours, the numerical optimization would take days or more likely weeks, and Bayesian methods would be entirely out of reach. This illustrates that our method enables fast and accurate parameter recovery which no previous methods are capable of.